# Differential phase contrast from electrons that cause inner shell ionization Michael Deimetry Timothy C. Petersen

### v1

@@ 4 Case study – thick specimen
The cross-section expression for electron energy loss spectroscopy is given by [40]:
where is the layer thickness, is an effective scattering potential for the inelastic transition and incorporating the detector geometry777For notational simplicity, we have absorbed into the prefactors that more conventionally appear in front of Eq.25., and this model neglects elastic scattering after the inelastic scattering event. The blue dotted lines, labeled , in Fig.8(b)-(d) show the Ti L1-edge signal calculated in this model, assuming a 35 eV energy window above threshold and a very large detector collection angle. The radius of the disk integration region used to evaluate Eq.24 was chosen to be that containing of the effective scattering potential since it gave the most favourable agreement with the results of Eq.25. Nevertheless, and while not identical, the similarity in the plots supports the channelling behaviour being the primary cause of the oscillations seen.
The green lines in Fig.8(b)-(d) were calculated using the transition potential model of Eq.23, assuming an energy filter 10 eV above the Ti L1-edge and a large detector. Since this quantity is the would-be-observed intensity on an EELS spectrum at the corresponding energy loss, the signal is labelled “total EELS”. By including absorption due to thermal scattering after the ionization event and restricting to a single energy loss, the assumptions of this calculation differ somewhat from that of the cross-section expression. (Elastic scattering after the ionization event is also included, but that redistribution will not much affect the total inelastic intensity.) Nevertheless, the total EELS signal and the cross-section expression are in close agreement, suggesting those differences have only a small effect on the depth dependence of the signal.
@@
Figure 9: (a) Layer contributions (from a sub-set of layers) to the EFDPC images for the three different thickness and defocus parameter combinations indicated by the coloured boxes in Fig.8(c), specifically 78.1 Å (20 layers) and surface defocus (top), 156.2 Å (40 layers) and Å defocus (middle), and 234.3 Å (60 layers) and surface defocus (bottom), all assuming specimen, 300 keV beam electrons, 15.7 mrad aperture and energy filter placed 10 eV beyond the Ti -edge. A layer is taken to be the depth of a unitcell i.e. 3.9 Å. For each of the three cases respectively, (b), (c) and (d) plot the energy-filtered, integrated DPC (EFiDPC), integrated inelastic intensity (total EELS), scattering cross-section (, Eq.25), and the scattering cross-section with a top-hat function ( w/ top-hat, Eq.24) for the probe on the TiO column as a function of layer depth. Insets show the 2D EFiDPC maps about the TiO column. All plots are scaled such that maxima and minima align, with the units for the EFiDPC signal given on the left axis and those for the cross-section given on the right. The broad similarity between the plots results from the underlying channelling of the probe being common to each signal.
@@ 5 Conclusion
Within the phase object approximation applicable to thin samples, we have shown that the contrast is similar to that of elastic DPC imaging because the elastic DPC contrast is preserved within the momentum distribution of inelastically scattered electrons, over a spatial region corresponding to the range of appreciable ionization probability. For low energy losses and light elements, this means that the EFDPC signal at an energy loss corresponding to a particular element may include contrast from nearby atoms of different species, explaining why nitrogen atom sites are visible in the boron K-shell EFDPC experimental results of Haas and Koch [20]. However, for heavier elements the range of the ionization interaction is much narrower, and element selective DPC contrast is expected.
@@
EFDPC thus combines the advantages of STEM EELS and elastic DPC. Indeed, the strengths of each mitigate against limitations in the other. The delocalisation of EELS is mitigated by manifesting as a modulation of the gradient of the elastic potential. This may allow the position of atoms to be more precisely determined, for instance when probing orientation relationships between a thin surface layer and a substrate. Conversely, the limited thickness range of interpretable elastic DPC images has been mitigated by the incoherence of EELS, making direct interpretation possible in much thicker samples, extending the range of samples over which reliable DPC-like contrast can be achieved. Further work is needed in developing procedures to acquire core-loss filtered 4D STEM data to determine whether EFDPC might be convenient and/or competitive over separately recording EELS and elastic DPC images when seeking such structural insights. Irrespective, the present results offer some initial insights into the prospects of energy-filtered 4D STEM.
@@ Acknowledgments
This research is supported by an Australian Government Research Training Program Scholarship. This research was supported under the Discovery Projects funding scheme of the Australian Research Council (Project No.
FT190100619).

### v2

@@ 4 Case study – thick specimen
The cross-section expression for electron energy loss spectroscopy is given by [41]
where is the layer thickness, is an effective scattering potential for the inelastic transition and incorporating the detector geometry999For notational simplicity, we have absorbed into the prefactors that more conventionally appear in front of Eq.25., and this model neglects elastic scattering after the inelastic scattering event. The blue dotted lines, labeled , in Fig.9(b)-(d) show the titanium L1 edge signal calculated in this model, assuming a 35 eV energy window above threshold and a very large detector collection angle. The radius of the disk integration region used to evaluate Eq.24 was chosen to be that containing of the effective scattering potential since it gave the most favourable agreement with the results of Eq.25. Nevertheless, and while not identical, the similarity in the plots supports the channelling behaviour being the primary cause of the oscillations seen.
The green lines in Fig.9(b)-(d) were calculated using the transition potential model of Eq.23, assuming an energy filter 10 eV above the titanium L1 edge and a large detector. Since this quantity is the would-be-observed intensity on an EELS spectrum at the corresponding energy loss, the signal is labelled “total EELS”. By including absorption due to thermal scattering after the ionization event and restricting to a single energy loss, the assumptions of this calculation differ somewhat from that of the cross-section expression. (Elastic scattering after the ionization event is also included, but that redistribution will not much affect the total inelastic intensity.) Nevertheless, the total EELS signal and the cross-section expression are in close agreement, suggesting those differences have only a small effect on the depth dependence of the signal.
@@
5 Practical considerations
Having focused primarily on qualitative interpretation, contrast formation mechanisms and imaging dynamics, we have thus far presented results assuming infinite dose. Though the experimental results of Haas and Koch [20] clearly show EFDPC to be achievable even from a monolayer, one challenge for this imaging mode is that the cross-section for core-loss scattering is typically several orders of magnitude lower than that of elastic scattering, meaning high doses will be required to obtain sufficient signal to be discernable above shot noise. But before considering typical doses required for EFDPC imaging, a comment on units is warranted.
We display EFDPC images simulated using the transition potential approach with arbitrary units not because the units are unknown but because they are not intuitive. There are two reasons for this. First, by considering ionization to continuum states the transition potential calculation of core-loss filtered diffraction intensities via Eq. (2) has units of eV-1, which follows from the normalisation of Eq. (6) in the Supplementary Material. Putting core-loss-filtered diffraction intensities on the same scale as the wavefield normalisation requires integration over a range of energy losses, with separate inelastic scattering calculations required at each energy loss, a complication that was unnecessary for exploring the contrast formation mechanisms. Second, the scale of the first moment evaluation can depend on the scale of the wavefield. In practice, elastic DPC is usually evaluated not by Eq. (12) but rather by the normalised form
eliminating any wavefunction normalisation to leave the result in units of the diffraction plane coordinates (often converted to a scattering angle). This works because the total recorded intensity (the denominator in Eq. (26)) depends only weakly on probe position.101010That it depends on probe position at all results from scattering outside the finite detector extent when the probe is above strongly scattering columns, and the normalisation of Eq. (26) can to some extent compensate for this. Core-loss-filtered 4D STEM is very different: the total inelastic signal varies considerably depending on where the probe is relative to columns of the element selected. Simulations (not shown) indicate that a normalisation like Eq. (26) for EFDPC reduces the qualitative similarity with elastic DPC images. But the consequence of not performing such a normalisation is that the first moment evaluation of Eq. (12) is significantly affected by the scale of the wavefunction, preventing intuitive display of EFDPC images as maps of the scattering angle.
Despite this, we can still explore the typical dose required for EFDPC imaging. Figure10 shows simulated elastic DPC (left) and boron K edge EFDPC (right, assuming a 20 eV energy window above the ionization threshold111111Modelled using numerical integration over transition-potential-based simulations sampled every 2 eV in energy loss.) images for a monolayer hBN including realistic (Poisson-distributed) shot noise based on the number of electrons reaching the 4D STEM detector. Figure10(a) and (b) assumes the number of electron per probe position, , to be (a) (or e-/Å2 at the sampling shown) and (b) (or e-/Å2 at the sampling shown), chosen to be the lowest values, to the nearest order of magnitude, for which signal is clearly discernible above the noise. Thus in this case the EFDPC image requires times more dose than elastic DPC for comparable contrast. Whereas Fig.10(a) and (b) (indeed all simulations thus far) assumes perfect source coherence, Fig.10(c) and (d) further accounts for partial spatial incoherence, included here assuming a Gaussian effective source distribution with a standard deviation of 0.3 Å that broadens the features while approximately halving the signal. Figure10(c) and (d) use the same dose as (a) and (b), respectively. The lower contrast is seen to manifest as more pronounced noise but the signal remains visible, suggesting that realistic effective source distributions have only a modest effect on the dose required.
Figure 10: (a) Elastic DPC and (b) EFDPC (with a 20 eV energy window above the boron K edge) simulated images of the first moment along the -direction for a detector spanning around mrad in the - and -directions for monolayer hBN. Calculations assume (a) and (b) electrons per probe position (corresponding to doses of (a) e-/Å2 and (b) e-/Å2). (c,d) As per (a,b) but including partial spatial incoherence assuming a Gaussian effective source distribution with standard deviation 0.3 Å.
Figure11 compares elastic (left column) and inelastic (right column; with a 20 eV energy window above the titanium L1 edge) DPC image simulations for SrTiO3 at various thicknesses. (We revert to assuming assuming perfect coherence to enable direct comparison with Fig.8(a) and (c).) Again, the dose assumed for each case is, to the nearest order of magnitude, the lowest for which appreciable contrast is visible above the noise. Figure11(a) and (b) assumes a monolayer of SrTiO3, which, while very hard to achieve in practice, enables comparison with the monolayer hBN results in Fig.10(a) and (b). Cross-sections for ionizing different elements and shells will generally differ [43]. However, as per Eq.19, EFDPC images in the phase object approximation scale with both the ionization cross-section and the (gradient of the) elastic potential. Thus while the titanium L1 shell has a lower ionization cross-section than the boron K shell, Ti having a greater elastic potential than B results in the minimum dose needed for monoloayer SrTiO3 being lower than that for monolayer hBN.
Subsequent rows in Fig.11 show results for larger thicknesses of SrTiO3. Again the dose assumed for each case is, to the nearest order of magnitude, the lowest for which appreciable contrast is visible above the noise. To a thickness of around 4 nm, increasing the thickness reduces the dose required to achieve a comparable level of noise. Beyond that, the required dose is approximately constant, though the apparent noise level increases slightly with increasing thickness. This is broadly consistent with the thickness dependence of the signal as seen in Fig.8(a) and (c), where the signal initially grows, then saturates and finally fades somewhat with increasing thickness. Interestingly, for the first 4 nm the relative reduction in necessary dose with increasing thickness is greater for EFDPC than for elastic DPC. This is broadly consistent with the results of Fig.7, where in the phase object approximation the elastic signal increases linearly with thickness whereas the EFDPC signal increases quadratically with thickness (due to both the accumulation of projected potential and the increase of the number of atoms ionized). We attribute the necessary dose not exactly following that scaling to the noise depending not just on the number of electrons but also on the details of the diffraction intensity distribution [44]. One way to further increase the signal-to-noise would be to use a larger energy window. Another, if the sample is periodic, would be repeat unit averaging.
Figure 11:
Elastic DPC (left column) and EFDPC (right column, with a 20 eV energy window above the titanium L1 edge) simulated images of the first moment along the -direction for a detector with an angular range extending out to 35 mrad for for thicknesses (a,b) 3.9 Å (1 layer), (c,d) 19.5 Å (5 layers), (e,f) 39.1 Å (10 layers), (g,h) 78.1 Å (20 layers), (i,j) 156.2 Å (40 layers) and (k,l) 234.3 Å (60 layers). The values above each image correspond to the number of electrons per probe position. For example, (a) and (b) respectively assume and electrons per probe position (corresponding to doses of e-/Å2 and e-/Å2, respectively). All calculations assume defocus to the midplane.
Isolating the core-loss contribution experimentally will require reliable background subtraction. Having identified the preservation of elastic contrast as the dominant mechanism leading to EFDPC contrast from core-loss events, it seems likely that some form of DPC contrast is also present in the background contribution. The approach of Haas and Koch [20] using - imaging has the advantage of providing spectra from which to fit the background, but data would need to be acquired sequentially at multiple slit orientations to adequately sample the momentum distribution. Conversely, the approach of core-loss-filtered 4D STEM as depicted in Fig.1 has the advantage of providing a detailed momentum distribution, but data would need to be acquired sequentially over multiple energy loss windows to effect background subtraction. Further experimental work is thus needed to determine the most effective approach to acquiring such data in practice.
@@ 6 Conclusion
Within the phase object approximation applicable to thin samples, we have shown that the contrast is similar to that of elastic DPC imaging because the elastic DPC contrast is preserved within the momentum distribution of inelastically scattered electrons, over a spatial region corresponding to the range of appreciable ionization probability. For low energy losses and light elements, this means that the EFDPC signal at an energy loss corresponding to a particular element may include contrast from nearby atoms of different species, explaining why nitrogen atom sites are visible in the boron K shell EFDPC experimental results of Haas and Koch [20]. However, for heavier elements the range of the ionization interaction is much narrower, and element selective DPC contrast is expected.
@@
EFDPC thus combines the advantages of STEM EELS and elastic DPC. Indeed, the strengths of each mitigate against limitations in the other. The delocalisation of EELS is mitigated by manifesting as a modulation of the gradient of the elastic potential. This may allow the position of atoms to be more precisely determined, for instance when probing orientation relationships between a thin surface layer and a substrate. Conversely, the limited thickness range of interpretable elastic DPC images has been mitigated by the incoherence of EELS, making direct interpretation possible in much thicker samples, extending the range of samples over which reliably interpretable DPC-like contrast can be achieved. Because ionization cross-sections are relatively low, in practice a very high dose is needed for sufficient counting statistics to discern differential phase contrast from core-loss scattered electrons. However, EFDPC depending on both the number of atoms ionized and the (gradient of the) total projected potential means that the signal strength initially scales quadratically with thickness (until saturating beyond about 4 nm), leading to some enhancement of signal-to-noise ratio that makes inelastic DPC imaging more achievable for thicker samples. Further work is needed in developing procedures to acquire core-loss filtered 4D STEM data to determine whether EFDPC might be convenient and/or competitive over separately recording EELS and elastic DPC images when seeking such structural insights. Irrespective, the present results offer some initial insights into the prospects of energy-filtered 4D STEM.
@@ Acknowledgments
This research is supported by an Australian Government Research Training Program Scholarship. This research was supported under the Discovery Projects funding scheme of the Australian Research Council (Project No. FT190100619).